A card is selected from a deck of 52 cards. Let A = Event that a...

  1. A card is selected from a deck of 52 cards.

Let A = Event that a diamond is selected

Let B = Event that a king is selected

            Find:   1) P(A), 2) P(B), 3) P(A&B), 4) P(A or B)
Answer #1

P(A) = 13/52 = 1/4

P(B) = 4/52 = 1/13

P(A&B) = 1/52

P(A or B) = P(A) + P(B) - P(A&B) = 1/4 + 1/13 - 1/52 = 4/13

press like